OpenStudy (anonymous):

solve for x : 3/2x-15=-12

OpenStudy (anonymous):

(3-30x)/2x = -12 3-30x = -24x -6x = -3 x = -1/2

OpenStudy (anonymous):

x = 1/2

OpenStudy (anonymous):

x = 1/2

OpenStudy (anonymous):

x=1/2

jhonyy9 (jhonyy9):

3x-30=-24 3x=6 x=6/3 x=2